# Coupled cluster method tailored with quantum computing

### v1

@@ III.3 Application to Cope rearrangement
For the chair and hexadiene ($C_{i}$ symmetry) geometries, we used the optimized structures from Ref. [72]; they used the analytic MR-CISD and multireference averaged quadratic coupled-cluster (MR-AQCC) gradient methods [73] to optimize the structures with a reference space of CAS(6o,6e) in a 6-31G*** basis set.
@@
Table 1: Activation energies in kcal/mol for 1,5-hexadiene Cope rearrangement calculated in the 6-31G*** basis set. The VQE calculations used the disentangled UCCSD ansatz and an active space consisting of 6 orbitals and 6 electrons [i.e., (6o, 6e)]. The experimental values were obtained from the computational study in Ref. [74], which was based on the experimentally measured enthalpy from Ref. [75]
@@ IV Conclusion
Nevertheless, for more complex systems or in the presence of real device errors such as depolarising noise, further verification is required to determine the extent to which QC-CBT-TCC works well. Although, in this study, we utilized the CBT method to approximate the wave function on the quantum computer, it is possible to explore extensions that combine TCC with methods that are more resilient to noise and statistical errors, such as quantum selected configuration interaction [76, 77].
@@ Appendix A Number of shot dependency of QC-CBT-T for LiH
In Figures 8 (a) and 8 (b) the results for LiH show a behavior similar to that of H2O and N2 in that increasing the number of measurements reduces the corrected energy distribution. This behavior can be explained by the higher accuracy and decreased uncertainty in determining the coefficients of a state on a quantum computer. Consequently, the obtained CBT state is more consistent, and therefore the expected energy can be determined with a higher degree of confidence.
Figure 8: Energy insecurity of QC-CBT-TCC(c) for LiH with a different number of measurements for CBT. $N_{\text{sample}},N_{U}$, and $N_{V}$ were all measured in the specified quantities in the figure as the X-axis values. CBT was performed 1000 times for each setting to obtain an estimate for the statistical errors. (a) Number of measurements dependency at an equilibrium distance of 3.016 Bohr. (b) Number of measurements dependency at 6.032 Bohr, twice the equilibrium distance.

### v2

@@ III.3 Application to Cope rearrangement
For the chair and hexadiene ($C_{i}$ symmetry) geometries, we used the optimized structures from Ref. [82]; they used the analytic MR-CISD and multireference averaged quadratic coupled-cluster (MR-AQCC) gradient methods [83] to optimize the structures with a reference space of CAS(6o,6e) in a 6-31G* basis set.
@@
Table 1: Activation energies in kcal/mol for 1,5-hexadiene Cope rearrangement calculated in the 6-31G* basis set. The VQE calculations used the disentangled UCCSD ansatz and an active space consisting of 6 orbitals and 6 electrons [i.e., (6o, 6e)]. The experimental values were obtained from the computational study in Ref. [84], which was based on the experimentally measured enthalpy from Ref. [85]
@@ IV Conclusions
Nevertheless, for more complex systems or in the presence of real device errors such as depolarising noise, further verification is required to determine the extent to which QC-CBT-TCC works well. Although, in this study, we utilized the CBT method to approximate the wavefunction on the quantum computer, it is possible to explore extensions that combine TCC with methods that are more resilient to noise and statistical errors, such as quantum selected configuration interaction [86, 87]. Another potential avenue for further development of this method is to create a self-consistent version [88, 89] that iterates between the calculation of the active space wavefunction and the optimization of the coupled cluster amplitudes in the outside of the active space, although such a self-consistent approach multiplies the computational cost associated with the method.
Finally, the performance of a tomography method depends heavily on the number of shots available. Matchgate/Fermionic shadows has good asymptotic scaling and can be highly efficient if enough shots are available [90, 91]. On the other hand, these shadow tomography methods may not always be the best approach given realistic shot budgets [44]. We chose CBT for this study because it requires shallower circuits and can be effective with small shot budgets by adjusting the $R$ parameter. Nevertheless, it is an open question as to which method is better for extracting CI coefficients with a limited number of shots.
@@ Appendix A Number of shot dependency of QC-CBT-T for LiH
In Figures 8 (a) and 8 (b) the results for LiH show a behavior similar to that of H2O and N2 in that increasing the number of measurements reduces the corrected energy distribution. This behavior can be explained by the higher accuracy and decreased uncertainty in determining the coefficients of a state on a quantum computer. Consequently, the obtained CBT state is more consistent, and therefore, the expected energy can be determined with a higher degree of confidence.
Figure 8: Energy insecurity of QC-CBT-TCC(c) for LiH with a different number of measurements for CBT. $N_{\text{sample}},N_{U}$, and $N_{V}$ were all measured in the specified quantities in the figure as the X-axis values. CBT was performed 1000 times for each setting to obtain an estimate for the statistical errors. The box indicates the range from the first to the third quartile called the inter-quartile range (IQR), with the median drawn in the middle. The box contains $50\%$ of the data. Whiskers are the lines extending 1.5 times the IQR from the first and third quantiles. Data points that exceed the whiskers are considered outliers (fliers) and are represented as single dots. (a) Number of measurements dependency at an equilibrium distance of 3.016 Bohr. (b) Number of measurements dependency at 6.032 Bohr, twice the equilibrium distance.